OpenStudy (anonymous):

Complete the square. -16t^2 - 8t = 10 Thanks!! :)

OpenStudy (astrophysics):

Start off by moving the 10 to the left side, to complete the square , say we have \[ax^2+bx+c = 0\] we do the following \[x^2+2 \frac{ a }{ 2 }x+ \left( \frac{ a }{ 2 } \right)^2-\left( \frac{ a }{ 2 } \right)^2\]

OpenStudy (anonymous):

Ok, what next?

OpenStudy (anonymous):

So far we have 16t2+8t+10=0 ?

OpenStudy (anonymous):

K I have to go...this is taking too long. Thanks

OpenStudy (anonymous):

Thanks but I just can't handle waiting for someone on a third-world internet connection

OpenStudy (anonymous):

I love explanations when they don't take 30 minutes